Decision

As part of their application for a loan to buy Lakeside Farm, a property they hope to develop as a bed-and-breakfast operation, the prospective owners have projected:

Monthly fixed cost (loan payment, taxes, insurance, maintenance)

$6000

Variable cost per occupied room per night

20

Revenue per occupied room per night

75

a. Write the expression for total cost per month. Assume 30 days per month.
b. Write the expression for total revenue per month.

c.

If there are 12 guest rooms available, can they break even? What percentage of rooms would need to be occupied, on average, to break even?

variables in a production process are those that cannot be controlled by the manager
Answer #1

The answer to Q a:

To derive the expression, first identify the fixed cost and the variable cost. The expression is Y = aX+ b. Here, b is fixed cost, i.e. 6000 and ‘a’ is a variable cost, i.e. 20*30. As the variable cost per occupied room is 20 per night, to get the monthly cost, it is multiplied by 30. Here, X is the number of rooms occupied.

Y(x) = 6000 + 20(30) X (monthly)

The answer to Q b:

Revenue per occupied room per night is 75, so the revenue expression will be as follows.

R(x) = 75(30) X (monthly)

The answer to Q c:

Fixed cost = 6000

Variable cost per room per night = 20

Revenue per room per night = 75

Breakeven point = Fixed cost / (revenue per room per night – Variable cost per room per night) *30= 6000 / (75 – 20) * 30 = 6000 / 55 * 30 = 6000 / 1650 = 3.64 = 4 rooms per night. The hotel has enough rooms to break even.

Occupancy rate = (Break-Even point / total rooms)*100 = (4/12) *100 = 100/3 = 33.33 %

Please upvote, if you find the answer helpful.